OpenStudy (anonymous):

if i have 1/3 and 1/7 of the circle unshaded whats the shaded part of the circle

OpenStudy (anonymous):

Um that's… Let's see if 1/3 isn't shaded than 2/3 is shaded If 1/7 isn't shaded than 6/7 is shaded

TheSmartOne (thesmartone):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\]

TheSmartOne (thesmartone):

It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

So your answer would be \[1-(\frac{ 1 }{ 3 }+\frac{ 1 }{ 7 })\]

TheSmartOne (thesmartone):

\[1-(\frac{ 7 }{21 }+\frac{ 3 }{ 21})=1-\frac{ 10 }{ 21 }\]

TheSmartOne (thesmartone):

\[\frac{ 21 }{ 21 }-\frac{ 10 }{ 21 }\]